Find the general term of a sequence whose first four terms are given.
28.)
2, -1,-4,-7,…

Answer by stanbon(48568)

Find the general term of a sequence whose first four terms are given.
2, -1,-4,-7,…
--------------
The sequence has a common difference of d=-3 so it is arithmetic.
--------------
Formula: a(n)=a(1)+(n-1)d
Your sequence formula:
a(n)=2+(n-1)(-3)
a(n)=2-3n+3
a(n)=5-3n
==============
Cheers,
Stan H.